** SHOW YOUR WORK !! *

.

STEM Inc. wants to expand their facility. They are considering proposals from two builders named “Builder A” and “Builder B”. STEM’s annual fixed costs for facilities from Builder A and Builder B are respectively $35,000 and $50,000. The variable cost for the product made in the facility from Builder A is $16.00 per unit and for the facility by Builder B, the variable cost is $10.00 per unit. The revenue generated by the product produced in the additional facility is $20 per unit, and the estimated annual output from the additional facility is 6000 units for the facility by each builder.

.

(a) On the basis of total cost, which builder should be preferred? Show your workings by determining the total cost for each builder.

(b) Determine the breakeven quantity for facility by Builder A.

(c) Determine the breakeven quantity for facility by Builder B.

(d) State how the results in (2b) and (2c) support the correct answer in (2a).
Answer #1

a) Profit = (Revenue - Expense)

Expense = Fixed Cost + Variable cost*Total Output quantity

Revenue = Total Output Quantity"Selling Price

For Builder A

Profit = (6000*20) - (35000 + 16*6000)

= 12000 - 131000

= -$11000

Hence the Builder A will incur a loss of $11000

Builder B

Profit = (6000*20) - (50000 + 10*6000)

= 120000 - 110000

= $10000

The builder used in this context would be builder B.

35000

b) let the break even quantity be X, then:

35000 + 16*X = 20*X

4*X = 35000

X = 8750

c) let the break even quantity be Y, then:

50000 + 10*X = 20*X

10*X = 50000

X = 5000

d) If we look at the breakeven points of the Builde A and B we will observe that Builder B has a lower break even point hence it will be more profitable as the range is lower that Builder A and this proves that the deduction in question (a) is true.